# Robot Drummer: Learning Rhythmic Skills for Humanoid Drumming

### v1

@@ Abstract
Humanoid robots have seen remarkable advances in dexterity, balance, and locomotion, yet their role in expressive domains, such as music performance, remains largely unexplored. Musical tasks, like drumming, present unique challenges, including split-second timing, rapid contacts, and multi-limb coordination over pieces lasting minutes. In this paper, we introduce Robot Drummer, a humanoid system capable of expressive, high-precision drumming across a diverse repertoire of songs. We formulate humanoid drumming as sequential fulfillment of timed-contacts and transform drum scores in to a Rhythmic Contact Chain. To handle the long-horizon nature of musical performance, we decompose each piece into fixed-length segments and train a single policy across all segments in parallel using reinforcement learning. Through extensive experiments on over thirty popular rock, metal, and jazz tracks, our results demonstrate that Robot Drummer consistently achieves high F1 scores. The learned behaviors exhibit emergent human-like drumming strategies, such as cross-arm strikes, and adaptive sticks assignments, demonstrating the potential of reinforcement learning to bring humanoid robots into the domain of creative musical performance. Project page: robot-drummer.github.io
@@ I Introduction
In this work, we aim to address this gap by introducing Robot Drummer, a system with humanoid robot that learns to play a variety of songs. The robots’ task is to perform a set of drumming patterns based on MIDI (Musical Instrument Digital Interface) transcription of sheet music, replicating the nuanced timing and expressivity typical of human drummers. This task requires our system to master three key challenges (i) temporal precision, ensuring each drum strike occurs at the exact moment required by the rhythm; (ii) spatial coordination, enabling the robot to strike multiple drums simultaneously or in quick succession; and (iii) dynamic adaptation, adjusting the motion in response to changes in tempo and musical intensity. While most of the current research in humanoid robotics has focused on loco-manipulation and whole-body control, this work aims to extend the frontier by endowing humanoid robots with creative and expressive skills.
@@ III Robot Drummer: Learning Rhythmic Skills
Our main goal is to enable a humanoid robot to perform a wide range of songs by striking correct sequence of drums (as illustrated in Figure 2), based on MIDI transcription of sheet music. Each drum is modeled as a collision body which registers an impact when struck with a sufficient force. The synthesizer then produces a corresponding sound upon detecting drum strike. Drumming is a highly dynamic and contact-rich task that requires executing a series of precisely timed strikes over a long temporal horizon. The length of songs can range from few minutes to well over 6 minutes. The long horizon nature of the task combined with the variety of rhythmic patterns, and reward sparsity in high dimensional space makes RL exploration particularly challenging. In this work, we tackle these challenges by introducing two main approaches: (1) formulating drumming as a sequential fulfillment of contacts, which we refer to as Rhythmic Contact Chain; (2) temporal decomposition of a long-horizon musical piece into multiple segments, each trained in parallel with a unified policy. This formulation provides a general framework for learning rhythmic skills through task-agnostic contact rewards, improving sample efficiency of exploring in high-dimensional space. An overview of our approach is illustrated in Figure 3.
@@ III-A Music Representation
While music can be represented as raw waveforms or spectrograms, these formats present challenges for robots, particularly when processing continuous data and designing appropriate reward functions. We use MIDI as a musical representation, which provides a standardized way to encode instruments’ timing in digital music. We begin with a MIDI file containing all the instruments of a musical piece. To isolate the percussion, we extract only the events on the dedicated drum channel, yielding a sequence of time-stamped “note-on” messages that correspond to drums hits. Each ”note-on” event contains pitch (MIDI note number), velocity, start time, and end time. From this drum-only track, we manually map each distinct MIDI note number to a physical drum in drum kit (see Figure 2). This mapping is song-specific and created after identifying the unique MIDI note pitches that are present in each song.
@@ V-C Generalist Policies
Combining multiple rhythmic patterns (e.g. irregular timing of Livin’ on a Prayer and Fire’s fast tempo) in one policy introduces negative transfer, impacting the policy’s ability to generalize. Several factors likely contribute to this generalization difficulty. The songs vary greatly in their spatial layout (e.g. 2-6 drums) and temporal structure, so the policy must accommodate conflicting demands. As our specialist results show, increasing the number of drums and temporal irregularity already correlate with lower performance. The generalist policy must cover both simple and complex patterns, which appears to impact it’s ability to perform well on a specific pattern. We observed that the generalist policies tend to adopt ’safer’ strategies, for instance, they often focus on repeatedly used drum targets and ignores others, much like the Roxanne’s specialist policy in Section V-A.
@@ VI Discussion
Our experiments highlighted both the strengths and limitations of RL in expressive humanoid drumming. Specialist policies mastered songs with high rhythmic fidelity, often with F1 scores above 0.9. The robot not only kept perfect timing during drumming performances but also executed complex drums fills and rolls (see Figures 4 and  6). Notably, the learned behaviors showed emergent spatial strategies: for example, the robot performed multiple cross-over strikes (hitting cymbal with the right stick and the left hand reaching for snare) while performing Livin’ on a Prayer. Such maneuvers were not explicitly encoded in the reward; they arose as the policy searched for efficient ways to reach distant drums.
@@ VII Conclusion
In this work, we have introduced Robot Drummer, a general RL-based system that enables a humanoid robot to perform complex drumming patterns. Our key contributions include modeling drumming as a rhythmic contact chain with a temporal segmentation approach and dense contact rewards that together allow training on long rhythmic sequences. The learned policies achieve near-perfect timing on many songs and exhibit human-like strategies (e.g., cross-overs strikes, fast drum rolls, and dynamic sticks assignments) without explicitly encoding them. Through quantitative experiments, we identified key factors affecting performance: songs with many drums or highly irregular timing are substantially harder to master. We also explored the possibility of learning a generalist drumming policy on multiple songs. While many specialist policies achieved near perfect performance on individual songs, a single generalist policy scored much lower, highlighting the challenge for generalization in rhythmic control. Despite the challenges for multi-task RL, the specialist policies demonstrated remarkably fluid and expressive drumming skills in simulation, suggesting that RL-based humanoid control can bridge into the creative domain. In future work, we will bridge the sim‑to‑real gap by deploying Robot Drummer on physical hardware and investigate improvisation and style adaptation, enabling new forms of musical collaboration between robots and human musicians.

### v2

@@ Abstract
Humanoid robots have seen remarkable advances in dexterity, balance, and locomotion, yet their role in expressive domains such as music performance remains largely unexplored. Musical tasks, like drumming, present unique challenges, including split-second timing, rapid contacts, and multi-limb coordination over performances lasting minutes. In this paper, we introduce Robot Drummer, a humanoid capable of expressive, high-precision drumming across a diverse repertoire of songs. We formulate humanoid drumming as sequential fulfillment of timed contacts and transform drum scores into a Rhythmic Contact Chain. To handle the long-horizon nature of musical performance, we decompose each piece into fixed-length segments and train a single policy across all segments in parallel using reinforcement learning. Through extensive experiments on over thirty popular rock, metal, and jazz tracks, our results demonstrate that Robot Drummer consistently achieves high F1 scores. The learned behaviors exhibit emergent human-like drumming strategies, such as cross-arm strikes, and adaptive stick assignments, demonstrating the potential of reinforcement learning to bring humanoid robots into the domain of creative musical performance. Project page: robotdrummer.github.io
@@ I Introduction
In this work, we aim to address this gap by introducing Robot Drummer, a system with a humanoid robot that learns to play a variety of songs. The robot’s task is to perform a set of drumming patterns based on MIDI (Musical Instrument Digital Interface) transcription of sheet music, replicating the nuanced timing and expressivity typical of human drummers. This task requires our system to master three key challenges (i) temporal precision, ensuring each drum strike occurs at the exact moment required by the rhythm; (ii) spatial coordination, enabling the robot to strike multiple drums simultaneously or in quick succession; and (iii) dynamic adaptation, adjusting the motion in response to changes in tempo and musical intensity. While most of the current research in humanoid robotics has focused on loco-manipulation and whole-body control, this work aims to extend the frontier by endowing humanoid robots with creative and expressive skills.
@@ III Robot Drummer: Learning Rhythmic Skills
Our main goal is to enable a humanoid robot to perform a wide range of songs by striking the correct sequence of drums (as illustrated in Figure 2), based on MIDI transcription of sheet music. Each drum is modeled as a collision body which registers an impact when struck with a sufficient force. The synthesizer then produces a corresponding sound upon detecting drum strike. Drumming is a highly dynamic and contact-rich task that requires executing a series of precisely timed strikes over a long temporal horizon. The length of songs can range from a few minutes to well over 6 minutes. The long-horizon nature of the task combined with the variety of rhythmic patterns, and reward sparsity in high dimensional space makes RL exploration particularly challenging. In this work, we tackle these challenges by introducing two main approaches: (1) formulating drumming as a sequential fulfillment of contacts, which we refer to as Rhythmic Contact Chain; (2) temporal decomposition of a long-horizon musical piece into multiple segments, each trained in parallel with a unified policy. This formulation provides a general framework for learning rhythmic skills through task-agnostic contact rewards, improving sample efficiency of exploring in high-dimensional space. An overview of our approach is illustrated in Figure 3.
@@ III-A Music Representation
While music can be represented as raw waveforms or spectrograms, these formats present challenges for robots, particularly when processing continuous data and designing appropriate reward functions. We use MIDI as a musical representation, which provides a standardized way to encode instruments’ timing in digital music. We begin with a MIDI file containing all the instruments of a musical piece. To isolate the percussion, we extract only the events on the dedicated drum channel, yielding a sequence of time-stamped “note-on” messages that correspond to drum hits. Each ”note-on” event contains pitch (MIDI note number), velocity, start time, and end time. From this drum-only track, we manually map each distinct MIDI note number to a physical drum in drum kit (see Figure 2). This mapping is song-specific and created after identifying the unique MIDI note pitches that are present in each song.
@@ V-C Generalist Policies
Combining multiple rhythmic patterns (e.g. irregular timing of Livin’ on a Prayer and Fire’s fast tempo) in one policy introduces negative transfer, impacting the policy’s ability to generalize. Several factors likely contribute to this generalization difficulty. The songs vary greatly in their spatial layout (e.g. 2-6 drums) and temporal structure, so the policy must accommodate conflicting demands. As our specialist results show, increasing the number of drums and temporal irregularity already correlate with lower performance. The generalist policy must cover both simple and complex patterns, which appears to impact its ability to perform well on a specific pattern. We observed that the generalist policies tend to adopt safer strategies, for instance, they often focus on repeatedly used drum targets and ignore others, much like the Roxanne’s specialist policy in Section V-A.
@@ VI Discussion
Our experiments highlighted both the strengths and limitations of RL in expressive humanoid drumming. Specialist policies mastered songs with high rhythmic fidelity, often with F1 scores above 0.9. The robot not only kept perfect timing during drumming performances but also executed complex drum fills and rolls (see Figures 4 and  6). Notably, the learned behaviors showed emergent spatial strategies: for example, the robot performed multiple cross-over strikes (hitting cymbal with the right stick and the left hand reaching for snare) while performing Livin’ on a Prayer. Such maneuvers were not explicitly encoded in the reward; they arose as the policy searched for efficient ways to reach distant drums.
@@ VII Conclusion
In this work, we have introduced Robot Drummer, a general RL-based system that enables a humanoid robot to perform complex drumming patterns. Our key contributions include modeling drumming as a rhythmic contact chain with a temporal segmentation approach and dense contact rewards that together allow training on long rhythmic sequences. The learned policies achieve near-perfect timing on many songs and exhibit human-like strategies (e.g., cross-over strikes, fast drum rolls, and dynamic stick assignments) without explicitly encoding them. Through quantitative experiments, we identified key factors affecting performance: songs with many drums or highly irregular timing are substantially harder to master. We also explored the possibility of learning a generalist drumming policy on multiple songs. While many specialist policies achieved near perfect performance on individual songs, a single generalist policy scored much lower, highlighting the challenge for generalization in rhythmic control. Despite the challenges for multi-task RL, the specialist policies demonstrated remarkably fluid and expressive drumming skills in simulation, suggesting that RL-based humanoid control can bridge into the creative domain. In future work, we will bridge the sim‑to‑real gap by deploying Robot Drummer on physical hardware and investigate improvisation and style adaptation, enabling new forms of musical collaboration between robots and human musicians.